OpenStudy (anonymous):

Find the height, in feet, of the rock after 5 seconds in the air. A rock is launched from a cannon. It’s height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. After 1 second, the rock is 129 feet in the air; after 2 seconds, it is 236 feet in the air.

OpenStudy (anonymous):

@JoKeR0331

OpenStudy (anonymous):

someone please help

OpenStudy (anonymous):

pleaseeeee

OpenStudy (anonymous):

@unknownunknown

OpenStudy (unknownunknown):

Do we have the angle of the cannon, is it firing directly upward?

OpenStudy (anonymous):

it dosent say @unknownunknown

OpenStudy (unknownunknown):

Can you do a screenshot of the question?

OpenStudy (anonymous):

i cant its a test

OpenStudy (anonymous):

do you want me to put every detail? @unknownunknown

OpenStudy (unknownunknown):

We will need every detail.. that's insufficient information as it is.

OpenStudy (anonymous):

Help please A rock is launched from a cannon. It's height, h(x), can be represented by a quadratic function in terms of time, x, in seconds.After one second, the rock is 129 feet in the air; after 2 seconds, it is 236 feet in the air. Find the height, in feet, of the rock after 5 seconds in the air.

OpenStudy (anonymous):

@unknownunknown all of it

OpenStudy (unknownunknown):

Alright well, I'm going to assume a lot, so I can't see if this will be definitely correct, but the question contains incorrect information to answer properly. So we'll just have to proceed with what we have.. Let me confirm the equation to make sure, one sec.

OpenStudy (unknownunknown):

insufficient information*

OpenStudy (anonymous):

so you cant solve it at all? @unknownunknown

OpenStudy (anonymous):

@unknownunknown ?????

OpenStudy (unknownunknown):

Is there a multiple choice or you must write the answer?

OpenStudy (anonymous):

must write the answer @unknownunknown

OpenStudy (unknownunknown):

So, if we let y equal the displacement, y = 9.8*t^2/2 + initial velocity * time So first of all, we'll need to solve for the initial velocity, can you rearrange the equation to solve for v? y=9.8*t^2/2 + v * t

OpenStudy (unknownunknown):

Actually since it's coming down, we should make 9.8*t^2/2 negative. So y = -9.8*t^2/2 + v*t

OpenStudy (anonymous):

i think its a large number for feet @unknownunknown

OpenStudy (unknownunknown):

I'd say the answer is approximately 527, somewhere around there.

OpenStudy (anonymous):

its not @unknownunknown

OpenStudy (unknownunknown):

And how do you know that?

OpenStudy (anonymous):

i just tried

OpenStudy (anonymous):

@unknownunknown

OpenStudy (unknownunknown):

Ok wait, let me work it out

OpenStudy (unknownunknown):

Have you covered gravity in class? g = 9.8?

OpenStudy (anonymous):

no @unknownunknown

OpenStudy (unknownunknown):

Okay so. Let's say we have a function.. f(t) = -11*t^2 + 140*t If we plug into this equation t=1, we have -11 + 140 = 129 which satisfies the first part Then if we plug t=2, we have -11*4 + 140*2 = 236 which satisfies the second part

OpenStudy (unknownunknown):

So then for f(5) = -11*25 + 140*5 = 425

OpenStudy (anonymous):

right @unknownunknown

OpenStudy (unknownunknown):

Okay, good.

OpenStudy (anonymous):

can you help with another one @unknownunknown

OpenStudy (anonymous):

Widget wonders produces widgets. They have found that the cost, c(x), of making x widgets is a quadratic function in terms of x. The company also discovered that it costs $23 to produce 2 widgets, $103 to produce 4 widgets, and $631 to produce 6 widgets.

OpenStudy (anonymous):

@unknownunknown